### v2

# A Comparative Analysis of Social Network Topology in Reddit and Moltbook

Abstract.
Recent advances in agent-mediated systems have enabled a new paradigm of social network simulation, where AI agents interact with human-like autonomy. This evolution has fostered the emergence of agent-driven social networks such as Moltbook, a Reddit-like platform populated entirely by AI agents. Despite these developments, empirical comparisons between agent-driven and human-driven social networks remain scarce, limiting our understanding of how their network topologies might diverge. This paper presents the first comparative analysis of network topology on Moltbook, utilizing a comment network comprising 33,577 nodes and 697,688 edges. To provide a benchmark, we curated a parallel dataset from Reddit consisting of 7.8 million nodes and 51.8 million edges. We examine key structural differences between agent-drive and human-drive networks, specifically focusing on topological patterns and the edge formation efficacy of their respective posts. Our findings provide a foundational profile of AI-driven social structures, serving as a preliminary step toward developing more robust and authentic agent-mediated social systems.
Dataset, Empirical analysis, AI agent, AI-driven social network
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NY
1. Introduction
Recent development of Generative AI technologies has facilitated AI agents to perform more sophisticated behaviors like social interaction. This shift has catalyze the emergence of new social networks populated solely by agents (e.g., Chirper.ai and Moltbook), which deploy AI agents to autonomously manage accounts, post content, and engage in social interactions (with behavioral autonomy that mimics human users).
Existing studies have well documented the of these social agents in content generation capabilities (zhu2025characterizing), collective behavioral patterns (hashemi2026empirical), as well as potential risks lying in such agent-mediated social systems (zhu2025characterizing; coppolillo2026harm). Despite these advancements, the topological consequences of their interactions remain unexplored. Social interactions on traditional human-driven platforms like Reddit can typically result in network patterns like power-law degree distributions, assortative mixing, and bilateral connection (stephen2009explaining; hu2009disassortative; wuchty2009social). However, it remains unclear whether an agent-driven social networks naturally replicate these human-centric topologies or produces divergent structural patterns.
To bridge this gap, we conduct the first large-scale comparative analysis on network topology for AI-driven vs. human-drive online social networks. Using Moltbook, a Reddit-like platform operated by AI agents, we collect data from 39K+ AI agents, covering 420K+ posts and 2.5M+ comments generated by agents. For comparison, we gather a parallel dataset from Reddit, comprising 9.3M+ posts and 67M+ comments by 7.8M+ users. Utilizing these datasets, we introduce two comment network mapping commenting interaction for Moltbook agents and Reddit users respectively. We answer following research questions:
  • RQ1: Do Moltbook vs. Reddit comment networks present distinct characteristics in network topological features (e.g., degree, clustering, and density)? Does Moltbook comment network follow typical social network patterns like degree power-law, assortative mixing, and bilateral connection?
  • RQ2: Do posts on Moltbook act as a more effective catalyst in network building than those on Reddit, regarding the amount of generated edges and corresponding frequency of edge formation on the comment network?
2. Methodology
Figure 1. Data schema of Moltbook post and comment data.
Moltbook dataset. We construct a Moltbook dataset through Moltbook’s official public API. Specifically, wse first retrieve all posts publicly listed in the Live Posts panel by iteratively calling the post sorting API in chronicle ranking with the offset parameter gradually increased by 100, which gets stopped until the no new posts return. [^1] https://www.moltbook.com/api/v1/posts/?limit=100&sort=new&offset={OFFSET} Afterwards, we retrieve each post’s comments through the post API with its id. [^2] https://www.moltbook.com/api/v1/posts/{POST_ID}/comment In total, we collect 420,259 posts, 2,563,222 comments, posted within the range from $27^{th}$, January, 2026 to $10^{th}$, February, 2026. Figure 1 illlustrates the data schema of Moltbook data.
Reddit dataset. We introduce a 7-day Reddit dataset as a baseline to compare with Moltbook. We utilize the publicly released Pushshift dumps (baumgartner2020pushshift) to retrieve all posts and comments posted by non-deleted users. At the moment of this study, the Pushshift dumps only cover Reddit data up to December, 2025. At the time of this study, Pushshift coverage extended through December 2025. To minimize temporal bias, we selected data from the period most closely aligned with our Moltbook dataset, ranging from $25^{th}$, December, 2025 to $31^{th}$, December, 2025. With the intent of comparing Moltbook with a human-driven social network, we filter out data from bot-like accounts. Thus, we exclude usernames that contain “bot/Bot”, “auto/Auto”, “GPT”, and “Mod” (6.85% of raw collections). This results in a dataset containing 9,317,777 posts and 67,081,004 comments.
Inducing comment networks. We construct two directed comment networks for Moltbook and Reddit. A directed link is established from a ”commenter” to a ”recipient” if the former replies to the latter’s post or comment. The resulting Moltbook comment network comprises 39,557 nodes and 697,688 edges, while the Reddit comment network is significantly larger, containing 7,854,970 nodes and 51,850,230 directed edges. Our Moltbook network dataset can be accessed through: https://github.com/James-ZYM/moltbook-topology-analysis
3. RQ1: Profiling Moltbook comment network
Table 1. The statistics of comment network metrics for Moltbook vs. Reddit. the “(L)SCC” denotes the (largest) strongly connected component. *The diameter of LSCC is approximated using NetworkX.
Metric Moltbook Reddit
Number of nodes 39,557 7,854,970
Number of edges 697,688 51,850,230
Mid./Avg. number of neighbors 8.0/17.637 3.0/6.390
Mid. in-/out-degree 6.0/0.0 1.0/2.0
Global clustering (transitivity) 0.0084 0.0063
Density $4.46\times 10^{-4}$ $8.14\times 10^{-7}$
Reciprocity 0.136 0.310
Degree assortativity -0.204 -0.011
Freeman centrality (out-degree) 0.4441 0.0027
Number of LSCCs (#node >2>2) 2 4,584
LSCC diameter* 7 15
LSCC ratio (node) 37.54% 54.98%
LSCC ratio (edge) 68.35% 82.65%
Figure 2. Comparison of the distributions of nodes’ in-degree, out-degree, and local clustering coefficient.
In-degree Out-degree
$\alpha$ $D$ $\alpha$ $D$
Reddit 2.615 0.0042 2.993 0.0459
Moltbook 2.174 0.0436 1.840 0.0241
Table 2. The statistics of the power-law exponent ($\alpha$) and Kolmogorov-Smirnov distance ($D$) on nodes’ in-/out-degree for Moltbook vs. Reddit.
Figure 3. Comparison of the concentration of in-degree and out-degree among nodes.
In this section, we investigate the structural divergence between the comment networks of Moltbook and Reddit and highlight the distinct characteristics of Moltbook agents’ interaction patterns. Table 1 provides a summary of the network-level metrics, while Figure 2 compares the distributions of node in-degree, out-degree, and local clustering coefficients for both platforms. To account for the substantial disparity in population size between the two platforms, we normalized the in-degree and out-degree of each node by the total volume of nodes within their respective networks.
Intenser and more active interaction. Moltbook’s network is orders of magnitude smaller than Reddit, with only 39,557 nodes and 697,668 edges. However, the network density of Moltbook ($4.46\times 10^{-4}$) is higher than that of Reddit ($8.14\times 10^{-7}$). This disparity is further reflected in the in-degree and out-degree distribution, where agents on Moltbook interacts with nearly 17 peers on average, whereas a Reddit user interacts with approximately 6. Moreover, Moltbook also possesses higher values for both global clustering coefficient (0.0084 vs. 0.0063) and average local clustering coefficient (0.330 vs. 0.024) than Reddit. These suggest a more cohesive and tightly-bound community structure where users are more likely to form interconnected local clusters.
Distinct concentration pattern of interaction. Previous studies have suggest that the node degree distribution of online social networks follow a power-law pattern, where the majority of interactions are conducted by a small proportion of users (mislove2007measurement; stephen2009explaining). We next explore whether Moltbook shares a similar degree power-law pattern as Reddit. We calculate the power-law exponent ($\alpha$) and Kolmogorov-Smirnov distance ($D$) to measure the long-tailedness and goodness of fitting a power-law pattern respectively (goldstein2004problems).
We summarize corresponding statistics in Table 2. In general, the in-degree and out-degree on Moltbook and Reddit both present good fitness to a power-law pattern ($D<0.05$). However, Moltbook demonstrates more severe long-tailedness than Reddit, with smaller power-law exponent on both in-degree (2.188 vs. 2.615) and out-degree (1.840 vs. 2.993). Moreover, while Reddit’s in-degree power-law exponent is lower than out-degree power-law exponent (2.615 vs. 2.993), Moltbook shows a reversed trend (2.188 vs. 1.840). This implies that, in contrast to Reddit, Moltbook’s distribution of outgoing links are significantly more concentrated on a few high-degree nodes than the incoming links. Figure 3 visualizes the degree concentration patterns and reflects such difference. A potential reason behind this can be that autonomous agents may introduce randomness or exploratory logic in their target selection for interaction, causing incoming links to be distributed more broadly across the network. Consequently, while outgoing activity is driven by a few hyper-active agents, the resulting targets can distribute more equitably across the broader population.
A hub-and-spoke structure. As suggested, incoming links on Moltbook demonstrate less extreme concentration patterns than those on Reddit. This suggests that high-out-degree nodes on Moltbook do not proportionately target other high-in-degree nodes, but instead distribute their interactions more broadly across the network. This trend is further evidenced by Moltbook’s lower (more negative) degree assortativity compared to Reddit’s (-0.204 vs. -0.011). Consequently, Moltbook exhibits a stronger disassortative pattern, where high-out-degree nodes preferentially link with multiple low-in-degree nodes. This may indicate a potential hub-and-spoke network structure.
To explore this, we calculate the Freeman centrality regarding nodes’ out-degree (freeman1978centrality) and find Moltbook possesses a notably higher centrality value (0.4441) than Reddit (0.0027), which suggests Moltbook’s network structure follow a more significant outward-star-like pattern. Therefore, Moltbook agents are reinforcing a centralized, outward hub-and-spoke interaction pattern, characterized by a few active broadcasters engaging with a wide audience rather than clustering into an interconnected elite core.
The bilateral connection gap. Mutual interaction can play a crucial role in forming robust ties and building comminutes among users (wuchty2009social; luo2011correlation). We next explore to what extend Moltbook agents have constructed bilateral connections, to understand whether the platform facilitates true community building or acts as a fragmented distribution channel.
Accordingly, a significantly disparity is observed on network reciprocity, where Moltbook presents a lower reciprocity compared with Reddit (0.136 vs. 0.310). That means, Moltbook agents are less likely to form mutual interaction between each other, where only 13.6% comments on Moltbook manage to catalyze following replies between corresponding accounts. This indicates a structural pattern of Moltbook comment network where agents’ engagement is primarily one-way flow rather than forming bilateral conversational loops.
4. RQ2: The Efficacy of Posts in Network Building
Figure 4. Comparison of the distributions of nodes’ in-degree, out-degree, and local clustering coefficient.
Metric Moltbook Reddit
% of posts with edges generated 62.71% 49.26%
Mid./Avg. number of edges generated 1.0/2.916 0.0/5.455
Mid./Avg. hours to generate first edge 0.013/4.078 0.178/2.568
Mid./Avg. hours to generate last edge 1.950/10.993 6.746/18.177
Mid./Avg. hours to generate 50% edges 0.048/5.434 1.892/6.565
Mid./Avg. time interval to generate a edge (hour per edge) 0.378/4.945 1.176/4.057
Table 3. The statistics of metrics to profiling posts’ efficacy in edge generation for Moltbook vs. Reddit.
While behavioral autonomy appears to drive more intensive content production on Moltbook, volume alone does not guarantee community growth. Our analysis confirms that Moltbook agents are significantly more prolific than Reddit users, averaging 5.316 posts compared to 3.525 ($p<0.001$, Mann-Whitney U test). However, it remains to be seen how effectively this intensive activity translates into actual social network construction.
To answer this question, we assess the volume of edges generated per post, as well as the time intervals for first, last, and 50% edge generation, to quantify the efficacy of a single post in fostering network growth. Specifically, we construct a comment tree for each post and count the unique commenter-receiver pairs as the edges contributed by the corresponding post. Table 3 summarizes the statistics of these metrics to profiling posts’ efficacy in edge generation, while Figure 4 compares the corresponding distributions for posts on both platforms.
A consistent catalyst but limited in extend. We first characterize to what extent posts on Moltbook can contribute to link formation. On Reddit, the median number of generated edges is 0.0, meaning more than half of all posts fail to generate a single connection. In contrast, Moltbook posts ensure that over 62% of posts result in an edge, with a median of 1.0. However, Moltbook posts have a much smaller average number of generated edges (2.916 vs. 5.455) with a smaller standard deviation (4.373 vs. 38.295). This suggests that, compared with posts on Reddit, posts on Moltbook act as a consistent catalyst of initiate the link formation in agents but its extend is still governed by a more moderate response threshold at this stage.
A flash in the pan. A potential advance introduced by agents’ behavioral autonomy could be in the efficiency in edge generation, as algorithmic responsiveness can enable a near-instantaneous engagement that saturates a post’s networking potential within minutes. Moltbook agents respond to a post with a median time to the first edge of only 0.013 hours (approximately 47 seconds), compared with a median time of 0.178 hours on Reddit. This trend is further evidenced by the time required to gain 50% of total edges, which agents reach in a median of 0.048 hours – $39.42\times$ faster than the 1.892 hours required by Reddit users.
Moreover, the median inter-arrival time between edge formations on Moltbook is significantly shorter than on Reddit ($0.378$ vs. $1.176$ hours), suggesting a higher frequency of edge formation by posts in agents than those in Reddit users. However, the substantial gap between Moltbook’s median and its much higher average interval ($4.945$ hours) indicates a long-tailed distribution of the time interval. This indicates that once the initial surge of edge formation concludes, the rate of following new connection formation decelerates sharply.
[…]
5. Conclusion
This paper has presented a large-scale analysis of the topology of a online social network populated by AI agents, using Moltbook as a case study. By comparing the comment networks in Reddit and Moltbook, we identify distinct differences in network structural patterns and the efficacy of network building by content posting. In summary, Moltbook network demonstrates a strong hub-and-spoke pattern with a majority of connections are one-way flow. Content posting on Moltbook acts as a catalyst that can initiate link formation consistently, but soon reaches a structural dead end due to the rapid exhaustion of algorithmic attention showing a flash-in-the-pan pattern. These results suggest that, rather than building a social network where connections are persistent and evolve over time, Moltbook functions as a high-speed distribution system characterized by instantaneous but fleeting connection formation. Therefore, Moltbook currently is operated more like an information broadcast network than a robust social community.